OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -7/8 and point (2, 3). A. y – 2 = - (7/8)(x + 3) B. y – 2 = (7/8)(x + 3) C. y – 3 = -(7/8)(x - 2) D. y – 3 = (7/8)(x - 3)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

plug the given info into point slope form \[\Large y - y_1 = m(x - x_1)\]

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

A @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you mixed up x and y

OpenStudy (anonymous):

so C? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large \Large y - y_1 = m(x - x_1)\] \[\Large \Large y - y_1 = -\frac{7}{8}(x - x_1)\] \[\Large \Large y - 3 = -\frac{7}{8}(x - 2)\] yep it's C

OpenStudy (anonymous):

okay thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np